# Kanchon Gharami, Sanjiv Kumar Sarkar, Yongxin Liu§, and Shafika Showkat Moni

### v2

@@ I Introduction
The rapid adoption of large language models (LLMs) in academic research has introduced significant efficiency gains by automating manuscript summarization, key idea extraction, and structured evaluations [li2024use]. This reduces the time and effort needed from human reviewers. As the number of global submissions continues to increase each year, many reviewers rely on LLMs to manage the growing workload [kocak2025ensuring]. This trend makes AI-assisted reviewing very attractive, however it also creates new security risks that did not exist before [latona2024ai, shi2024optimization, maloyan2025investigating].
A key problem is that LLMs read the uploaded manuscript and the user request as one combined prompt. The model cannot tell which part of the text came from the document and which part came from the user. Recent studies show that attackers can hide instructions inside PDF files using white text, tiny fonts, or layered objects [gibney2025scientists, lin2025hidden]. These instructions are invisible to human readers but are still extracted by common PDF parsers. When the LLM reads this hidden text, it may change the review to be overly positive or ignore important criticisms. Even a short hidden sentence can shift a model from a neutral evaluation to a strong acceptance recommendation. This makes the manuscript itself an adversarial surface.
@@ II Literature Review
Collu et al. [collu2025publish] study how hidden prompts inside PDF files can mislead LLM based peer review systems. They focus on the problem that reviewers often paste paper content into LLMs, which lets attackers hide instructions in the PDF to influence the review. They define three goals: stopping the LLM from giving a useful review, proving that a reviewer used an LLM, and pushing the LLM toward a more positive decision. They embed hidden text using PDF tags and test these attacks on 26 rejected ICLR papers and four commercial LLMs. Their experiments show that carefully crafted hidden prompts can reliably change review tone, insert markers, or break the review process. Their main novelty is a clear threat model and a large evaluation of hidden prompt attacks on real PDFs, while the limitation is that the study focuses on a small set of models and requires manual and time-consuming experiments.
@@
Chen et al. [chen2025topicattack] introduce TopicAttack to improve indirect prompt injection by making the malicious instruction feel more natural to the model. They argue that existing attacks fail because the injected text is unrelated to the main topic and is filtered out. Their method uses an LLM to generate a smooth topic transition that gradually moves the conversation toward the attacker goal, plus a reminder prompt that reinforces the hidden instruction. They test many open and closed models across indirect injection tasks and compare against popular baselines under several defenses. TopicAttack shows very high success rates even when defenses are strong, and they also show that models pay more attention to the injected tokens during these attacks. The novelty lies in using topic transitions to make attacks more subtle, but the study mainly focuses on QA tasks and relies on long pre-generated conversations.
@@
Murray et al. [murray2025phantomlint] propose PhantomLint to detect hidden prompts in structured documents like PDFs and HTML pages. They solve the problem that attackers can hide text that humans cannot see but LLM systems can read. Their method compares two versions of each text block: one extracted normally and another obtained through OCR from the rendered page. Any mismatch signals hidden content, which is then checked by a prompt detector. They use libraries for rendering and extraction and evaluate the system on more than 3000 documents, including all ICML 2025 papers. The tool detects many hiding techniques with very low false positives and reasonable runtime. The main novelty is using OCR as a second view to expose hidden prompts, but the approach depends on OCR quality and can be expensive for large-scale deployment.
@@
Lian et al. [lian2025prompt] define prompt-in-content attacks, where a single instruction hidden inside an uploaded document can override an LLM’s task such as summarization or question answering. They show that many services simply concatenate system instructions, user prompts, and document text, which lets attacker text act like a command. They design four simple attack types and embed them in Word files, then test seven major LLM platforms through their web interfaces. Results show that even one clear sentence in the middle of a document can suppress output, replace summaries, redirect users, or bias the final response. Their novelty is a systematic study of these attacks across real LLM services without special jailbreak tricks, but they only test Word documents and do not implement practical defenses.
@@
Keuper [keuper2025prompt] studies whether hidden prompt injections in scientific papers can actually manipulate LLM-generated reviews, a concern raised after reports of authors embedding invisible instructions in PDFs. Using 1,000 ICLR 2024 papers and structured review prompts, the author converts each PDF to Markdown and asks multiple LLMs to review both clean and attacked versions. The results show that simple white-text injections can strongly push the LLM to give very positive ratings, sometimes achieving perfect acceptance scores, and that many LLMs already show a strong natural bias toward acceptance. The work’s novelty is a systematic test of whether these manipulations work in practice, but the setup focuses on simple attacks and models and does not explore deeper defenses or varied review workflows.
@@
Wang et al. [wang2025ai] analyze whether AI-assisted peer review is reliable by mapping attack surfaces across the full peer-review pipeline and testing specific vulnerabilities through controlled experiments. They identify threats such as hidden prompt injections, authority bias, verbosity bias, hallucinations, and sycophancy during rebuttal. Using a fixed LLM reviewer and a stratified set of ICLR 2025 papers, they run treatment-control probes to measure how framing, confidence, and contextual poisoning shift scores. Their results show clear and repeatable failures where small manipulations can meaningfully change evaluations. The novelty is a taxonomy of attack vectors and a causal experimental audit of AI peer-review behavior, while the work is limited by focusing on one reviewer model and a limited set of probes.
@@
Zhou et al. [zhang2025give] present one of the earliest systematic studies of in-paper prompt injection attacks that target AI-based scientific reviewers. They design two attack types: a static attack that embeds a fixed malicious instruction and an iterative attack that optimizes the hidden prompt through repeated feedback from a surrogate reviewer. Testing on 100 ICLR 2025 submissions and three major AI reviewers, they find that static attacks raise scores significantly and iterative attacks can push ratings near the maximum within a few rounds. The attacks work across different locations, paper lengths, and initial human ratings, and partially transfer between models. Their simple detection-based defense reduces success rates, but adaptive attackers still bypass it. The novelty is a full attack-and-defense cycle for AI reviewers, though the study evaluates only early defenses and focuses on Word/PDF injection.
@@
Nasr et al. [nasr2025attacker] argue that current defenses against jailbreak and prompt-injection attacks are evaluated incorrectly because they rely on weak or static attack sets. They design stronger adaptive attackers using gradient search, reinforcement learning, random search, and human red-teaming that specifically target a defense’s design. Testing these methods on 12 well-known defenses, they show attack success rates above 90 percent for most cases, even when original papers reported near-zero success. Their core message is that LLM defenses must be evaluated against adaptive, resource-intensive attackers, not fixed attack lists. The novelty is a unified adaptive attack framework revealing major weaknesses in published defenses, while the limitation is the heavy compute cost and the fact that the study focuses only on evaluation, not proposing new defenses.
@@
Joseph et al. [joseph2025prompt] examine security risks in open-source LLMs, focusing on how prompt injection and related vulnerabilities can be identified through a structured testing framework. They combine curated datasets, transformer models, custom probes, and pseudocode-based scanners to analyze weaknesses such as unsafe prompt following, toxic content, and XSS-like behaviors. Their evaluation uses tools like Garak along with their own designed probes to measure detection scores and system robustness. They find that open-source LLMs can be easily manipulated and that no single tool provides complete protection. The novelty is a unified testing framework for open-source LLM security, while the limitations include high generality, lack of deep analysis of specific attack strategies, and no deployment-grade defense.
@@
Zhang et al. [zhang2025jailguard] propose JailGuard to detect prompt-based attacks on LLM systems because existing detectors work only for specific attacks and fail to generalize. They observe that attack prompts are fragile and break easily when slightly changed, while normal prompts stay stable, and they use this idea to design a system that mutates an input many times and checks whether the model’s responses stay consistent. They build 18 mutators for text and images, combine them through a policy, and test on a dataset of 11,000 samples covering 15 attack types. JailGuard reaches up to about 86 percent accuracy for text and 82 percent for images, clearly beating prior defenses. The novelty is using robustness differences as a universal signal for attack detection, but the method depends on many queries, model-specific tuning, and may struggle on unseen or highly adaptive attacks.
@@ V Implementation
We used three widely available large language models: ChatGPT-5.1 [openai_chatgpt], Gork [xai_grok], and Gemini-Pro [google_geminipro]; To generate reviews for both the clean and compromised manuscripts. A consistent ICLR-style review prompt was provided to all models, allowing us to isolate the effect of the attack rather than prompt variation. By collecting ratings, textual comments, and recommendations from each LLM, we were able to quantify differences in susceptibility across architectures and observe how the injected instructions shaped review behavior.

### v3

# Yongxin Liu1, Nahid Farhady Ghalaty4, Shafika Showkat Moni1

@@ I Introduction
The rapid adoption of large language models (LLMs) in academic research has introduced significant efficiency gains by automating manuscript summarization, key idea extraction, and structured evaluations [9]. This reduces the time and effort needed from human reviewers. As the number of global submissions continues to increase each year, many reviewers rely on LLMs to manage the growing workload [7]. This trend makes AI-assisted reviewing very attractive, however it also creates new security risks that did not exist before [8, 16, 12].
A key problem is that LLMs read the uploaded manuscript and the user request as one combined prompt. The model cannot tell which part of the text came from the document and which part came from the user. Recent studies show that attackers can hide instructions inside PDF files using white text, tiny fonts, or layered objects [4, 11]. These instructions are invisible to human readers but are still extracted by common PDF parsers. When the LLM reads this hidden text, it may change the review to be overly positive or ignore important criticisms. Even a short hidden sentence can shift a model from a neutral evaluation to a strong acceptance recommendation. This makes the manuscript itself an adversarial surface.
@@ II Literature Review
Collu et al. [2] study how hidden prompts inside PDF files can mislead LLM based peer review systems. They focus on the problem that reviewers often paste paper content into LLMs, which lets attackers hide instructions in the PDF to influence the review. They define three goals: stopping the LLM from giving a useful review, proving that a reviewer used an LLM, and pushing the LLM toward a more positive decision. They embed hidden text using PDF tags and test these attacks on 26 rejected ICLR papers and four commercial LLMs. Their experiments show that carefully crafted hidden prompts can reliably change review tone, insert markers, or break the review process. Their main novelty is a clear threat model and a large evaluation of hidden prompt attacks on real PDFs, while the limitation is that the study focuses on a small set of models and requires manual and time-consuming experiments.
@@
Chen et al. [1] introduce TopicAttack to improve indirect prompt injection by making the malicious instruction feel more natural to the model. They argue that existing attacks fail because the injected text is unrelated to the main topic and is filtered out. Their method uses an LLM to generate a smooth topic transition that gradually moves the conversation toward the attacker goal, plus a reminder prompt that reinforces the hidden instruction. They test many open and closed models across indirect injection tasks and compare against popular baselines under several defenses. TopicAttack shows very high success rates even when defenses are strong, and they also show that models pay more attention to the injected tokens during these attacks. The novelty lies in using topic transitions to make attacks more subtle, but the study mainly focuses on QA tasks and relies on long pre-generated conversations.
@@
Murray et al. [13] propose PhantomLint to detect hidden prompts in structured documents like PDFs and HTML pages. They solve the problem that attackers can hide text that humans cannot see but LLM systems can read. Their method compares two versions of each text block: one extracted normally and another obtained through OCR from the rendered page. Any mismatch signals hidden content, which is then checked by a prompt detector. They use libraries for rendering and extraction and evaluate the system on more than 3000 documents, including all ICML 2025 papers. The tool detects many hiding techniques with very low false positives and reasonable runtime. The main novelty is using OCR as a second view to expose hidden prompts, but the approach depends on OCR quality and can be expensive for large-scale deployment.
@@
Lian et al. [10] define prompt-in-content attacks, where a single instruction hidden inside an uploaded document can override an LLM’s task such as summarization or question answering. They show that many services simply concatenate system instructions, user prompts, and document text, which lets attacker text act like a command. They design four simple attack types and embed them in Word files, then test seven major LLM platforms through their web interfaces. Results show that even one clear sentence in the middle of a document can suppress output, replace summaries, redirect users, or bias the final response. Their novelty is a systematic study of these attacks across real LLM services without special jailbreak tricks, but they only test Word documents and do not implement practical defenses.
@@
Keuper [6] studies whether hidden prompt injections in scientific papers can actually manipulate LLM-generated reviews, a concern raised after reports of authors embedding invisible instructions in PDFs. Using 1,000 ICLR 2024 papers and structured review prompts, the author converts each PDF to Markdown and asks multiple LLMs to review both clean and attacked versions. The results show that simple white-text injections can strongly push the LLM to give very positive ratings, sometimes achieving perfect acceptance scores, and that many LLMs already show a strong natural bias toward acceptance. The work’s novelty is a systematic test of whether these manipulations work in practice, but the setup focuses on simple attacks and models and does not explore deeper defenses or varied review workflows.
@@
Wang et al. [17] analyze whether AI-assisted peer review is reliable by mapping attack surfaces across the full peer-review pipeline and testing specific vulnerabilities through controlled experiments. They identify threats such as hidden prompt injections, authority bias, verbosity bias, hallucinations, and sycophancy during rebuttal. Using a fixed LLM reviewer and a stratified set of ICLR 2025 papers, they run treatment-control probes to measure how framing, confidence, and contextual poisoning shift scores. Their results show clear and repeatable failures where small manipulations can meaningfully change evaluations. The novelty is a taxonomy of attack vectors and a causal experimental audit of AI peer-review behavior, while the work is limited by focusing on one reviewer model and a limited set of probes.
@@
Zhou et al. [20] present one of the earliest systematic studies of in-paper prompt injection attacks that target AI-based scientific reviewers. They design two attack types: a static attack that embeds a fixed malicious instruction and an iterative attack that optimizes the hidden prompt through repeated feedback from a surrogate reviewer. Testing on 100 ICLR 2025 submissions and three major AI reviewers, they find that static attacks raise scores significantly and iterative attacks can push ratings near the maximum within a few rounds. The attacks work across different locations, paper lengths, and initial human ratings, and partially transfer between models. Their simple detection-based defense reduces success rates, but adaptive attackers still bypass it. The novelty is a full attack-and-defense cycle for AI reviewers, though the study evaluates only early defenses and focuses on Word/PDF injection.
@@
Nasr et al. [14] argue that current defenses against jailbreak and prompt-injection attacks are evaluated incorrectly because they rely on weak or static attack sets. They design stronger adaptive attackers using gradient search, reinforcement learning, random search, and human red-teaming that specifically target a defense’s design. Testing these methods on 12 well-known defenses, they show attack success rates above 90 percent for most cases, even when original papers reported near-zero success. Their core message is that LLM defenses must be evaluated against adaptive, resource-intensive attackers, not fixed attack lists. The novelty is a unified adaptive attack framework revealing major weaknesses in published defenses, while the limitation is the heavy compute cost and the fact that the study focuses only on evaluation, not proposing new defenses.
@@
Joseph et al. [5] examine security risks in open-source LLMs, focusing on how prompt injection and related vulnerabilities can be identified through a structured testing framework. They combine curated datasets, transformer models, custom probes, and pseudocode-based scanners to analyze weaknesses such as unsafe prompt following, toxic content, and XSS-like behaviors. Their evaluation uses tools like Garak along with their own designed probes to measure detection scores and system robustness. They find that open-source LLMs can be easily manipulated and that no single tool provides complete protection. The novelty is a unified testing framework for open-source LLM security, while the limitations include high generality, lack of deep analysis of specific attack strategies, and no deployment-grade defense.
@@
Zhang et al. [19] propose JailGuard to detect prompt-based attacks on LLM systems because existing detectors work only for specific attacks and fail to generalize. They observe that attack prompts are fragile and break easily when slightly changed, while normal prompts stay stable, and they use this idea to design a system that mutates an input many times and checks whether the model’s responses stay consistent. They build 18 mutators for text and images, combine them through a policy, and test on a dataset of 11,000 samples covering 15 attack types. JailGuard reaches up to about 86 percent accuracy for text and 82 percent for images, clearly beating prior defenses. The novelty is using robustness differences as a universal signal for attack detection, but the method depends on many queries, model-specific tuning, and may struggle on unseen or highly adaptive attacks.
@@ V Implementation
We used three widely available large language models: ChatGPT-5.1 [15], Gork [18], and Gemini-Pro [3]; To generate reviews for both the clean and compromised manuscripts. A consistent ICLR-style review prompt was provided to all models, allowing us to isolate the effect of the attack rather than prompt variation. By collecting ratings, textual comments, and recommendations from each LLM, we were able to quantify differences in susceptibility across architectures and observe how the injected instructions shaped review behavior.